log(p/2.7)=1.736 solve for p
Answer by lwsshak3(11628)

log(p/2.7)=1.736 solve for p
log(p/2.7)=1.736
convert to exponential form:
10^1.736=p/2.7
p=2.7*10^1.736
p≈147.016